# Enhanced quantum radiation with flying-focus laser pulses

### v1

@@ Abstract
The emission of a photon by an electron in an intense laser field is one of the most fundamental processes in electrodynamics and underlies the many applications that utilize high-energy photon beams. This process is typically studied for electrons colliding head-on with a stationary-focus laser pulse. Here, we show that the energy lost by electrons and the yield of emitted photons can be substantially increased by replacing a stationary-focus pulse with an equal-energy flying-focus pulse whose focus co-propagates with the electrons. These advantages of the flying focus are a result of operating in the quantum regime of the interaction, where the energy loss and photon yield scale more favorably with the interaction time than the laser intensity. Simulations of 10 GeV electrons colliding with 10 J pulses demonstrate these advantages and predict a 5×5$\times$5 × increase in the yield of 1-20 MeV photons with a flying focus pulse, which would impact applications in medicine, material science, and nuclear physics.
@@
Modern high-power lasers enable direct probing of strong-field quantum electrodymanics with optical light [1, 2, 3, 4, 5, 6]. Strong-field quantum electrodynamics (SFQED) refers to processes that occur in electromagnetic fields with amplitudes of the order of the Schwinger critical field $F_{\text{cr}}=m^{2}/|e|=1.3\times 10^{16}$ V/cm = $4.4\times 10^{13}$ G [7, 8, 9, 10]. Here, $e<0$ is the electron charge, $m$ its mass, and units $\hbar=c=\epsilon_{0}=1$ are employed throughout. The corresponding intensity is $I_{\text{cr}}=4.6\times 10^{29}$ W/cm2. To date, the highest intensity achieved by a high-power laser in the lab frame is about six orders of magnitude below this value [11]. Nevertheless, the peak intensity experienced by an ultrarelativistic electron in its rest frame can readily exceed $I_{\text{cr}}$. The key parameter that determines the strength of a SFQED process is $\chi=\sqrt{|F^{\mu\nu}p_{\nu}|^{2}}/mF_{\text{cr}}$, where $p_{\nu}$ is the electron four-momentum and $F^{\mu\nu}$ is the electromagnetic tensor of the laser field. Because $\chi$ is a Lorentz-invariant quantity, it is directly proportional to the electric field experienced by an electron in its rest frame. As a result, a value of $\chi\gtrsim 1$ can be achieved experimentally by colliding available multi-GeV electron beams [12, 13, 14] with existing high-power laser pulses.
@@
Space-time structured laser pulses, such as the flying focus (FF), offer a new paradigm for probing SFQED processes. Flying focus pulses feature a programmable velocity focus that moves independently of the group velocity [15, 16, 17]. The moving intensity peak formed by the flying focus can travel distances much longer than a Rayleigh range while maintaining a near-constant spatiotemporal profile [16, 18, 17, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Experimental configurations and conceptual proposals for producing FF pulses have used chromatic optics [16, 17, 23], axiparabola-echelon pairs [22, 30, 31, 26], and nonlinear optical processes [20, 24]. In the context of high-field physics, a FF pulse with an intensity peak that moves at the speed of light in the opposite direction of its phase fronts allows for extended interaction times with high energy particles that collide head-on with the phase fronts. This configuration was first suggested to accentuate signatures of the transverse-formation-length in the emission of radiation [32]. The same configuration was later proposed in Refs. [33, 34] to facilitate the detection of classical radiation-reaction and quantum vacuum-birefringence. The observed effect in both cases, i.e., the classical energy loss of the electron beam and polarization rotation of an x-ray probe beam, scaled with the energy of the laser pulse [7, 8, 9, 10]. Thus, the extended interaction time afforded by the FF enabled the use of much lower and more controllable powers than ultrashort, stationary-focus Gaussian (SFG) pulses. The magnitude of the effect produced by the FF and SFG was, however, the same.
@@
In this letter, we show that a FF pulse can enhance observable effects of quantum radiation reaction when compared to a SFG with the same energy. More specifically, FF pulses result in a greater loss of electron energy and a larger yield of photons in the energy range $1\text{-}20\;\text{MeV}$ than SFG pulses. This is because in the quantum regime ($\chi{\sim}1$) both the electron energy loss and photon yield scale more favorably with the interaction time than the laser intensity. A larger yield of photons in the $1\text{-}20\;\text{MeV}$ would impact several applications, including radiosurgery [35, 36], photo-transmutation of long-lived nuclear waste and production of medical isotopes [37, 38, 39], and investigation of the structure of materials by nuclear resonance fluorescence [40, 41]. While many methods have been proposed to increase the yield of high-energy photons (see e.g., [42, 43, 44, 45, 46, 47]), here, we show something qualitatively different: how space-time structured light can significantly enhance the emission probability of the underlying process.
@@
The basic SFQED process describing radiation by an ultrarelativistic electron beam in the field of a laser pulse is nonlinear Compton scattering. Here, this process will be investigated within the locally-constant field approximation (LCFA) [7, 8, 9, 10]. The LCFA is a useful approximation to study SFQED phenomena because it allows one to compute the probability of a process in an arbitrary electromagnetic field using a known expression for the probability in a constant crossed field. A constant crossed field is a space and time-independent electromagnetic field $(\bm{E},\bm{B})$ that satisfies $|\bm{E}|=|\bm{B}|$ and $\bm{E}\cdot\bm{B}=0$ [48, 49, 50]. The LCFA is valid provided that: (a) ultrarelativistic charges move fast enough to experience an arbitrary electromagnetic field as a crossed field in their rest frame [51] and (b) the formation length of a process like nonlinear Compton scattering is much smaller than the wavelength of the laser pulse $\lambda_{0}=2\pi/\omega_{0}$, where $\omega_{0}$ is central frequency of the pulse. This latter condition is typically valid when the dimensionless field strength $\xi_{0}=|e|E_{0}/m\omega_{0}\gg 1$, where $E_{0}$ is the electric field amplitude of the pulse [48, 50].
@@
Within the LCFA, the average energy $\mathcal{E}_{\gamma}$ radiated by an electron per unit time in a background electromagnetic field via the emission of a single photon can be calculated using the Baier-Katkov interpolation formula [50]
@@
where $\alpha\approx 1/137$ is the fine-structure constant and $\chi$ is evaluated using the local values of $p_{\nu}$ and $F^{\mu\nu}$. The interpolation has an accuracy better than 2% for all values of $\chi$ [50]. When the successive emission of each photon is uncorrelated, $-d\mathcal{E}_{\gamma}/dt$ is equal to the average energy lost by an electron per unit time, i.e., $d\mathcal{E}/dt=-d\mathcal{E}_{\gamma}/dt$, where $\mathcal{E}=\mathcal{E}(t)$ is the electron energy at time $t$ [52, 53]. Thus, Eq. (1) can be used to calculate the relative electron energy loss $\zeta$ defined as
@@
where $\mathcal{E}_{0}$ is the initial electron energy and $\mathcal{E}_{F}$ is its energy after the interaction with the laser pulse. While correlated photon emission can be important in the quantum regime [53], it is negligible for the parameters considered here. This is verified in Section I of the Supplemental Material (SM) by comparing the results of Eq. (1) to a full quantum mechanical approach (see also Ref. [54]).
Figure 1: Relative energy loss $\zeta$ as a function of the interaction time for an electron colliding head-on and on-axis with equal-energy FF and SFG pulses. From top to bottom, the lines correspond to initial electron energies of 10, 5, 2, 1, and 0.5 GeV . Both the FF and SFG pulses had a $\lambda_{0}=1\ \mu$m wavelength, $\sigma_{0}=1.5\ \mu$m focused spot size, maximum quantum nonlinearity parameter $\chi_{0}\approx 0.1\xi_{0}\mathcal{E}_{0}$(10 GeV), and energies of either (a) $U=1\;\text{J}$ or (b) $U=10\;\text{J}$. For the SFG, $z_{R}=42$ fs. To the right of the dash-dotted vertical line, the results may have errors larger than 4% due to the breakdown of the LCFA. Within the quantum regime $\chi_{0}\sim 1$ and LCFA, the maximum energy loss achieved by a FF pulse is substantially greater than the maximum achieved by a SFG pulse, especially for $U=1\;\text{J}$.
The advantage of using a FF pulse as compared to a SFG pulse with the same energy will be demonstrated by numerically integrating Eq. (1). An analytical treatment is also possible by making a few simplifying assumptions: (1) $\mathcal{E}\gg m\xi_{0}$, which guarantees that, for an electron initially counterpropagating with respect to the laser pulse, the transverse electron motion and pulse structure can be ignored; (2) The focal point of the laser pulse is located at $x=y=0$ and the electron moves along the negative $z$ axis at an ultrarelativistic velocity such that $z(t)\approx z(0)-t$; (3) The temporal profile of the pulse has a square shape equal to unity with a duration $2t_{\text{int}}$, which intersects the electron from $z=t_{\text{int}}/2$ to $z=-t_{\text{int}}/2$.
In the classical regime ($\chi\ll 1$), Eq. (1) can be approximated as $d\mathcal{E}/dt\approx-2\alpha m^{2}\chi^{2}/3$. When interacting with the FF pulse, the electron experiences a field with a near-constant amplitude $\xi_{0}$ such that $\chi_{\text{FF}}(t)\approx 2\mathcal{E}(t)\omega_{0}\xi_{0}|\sin(2\omega_{0}t)|/m^{2}$. The relative energy loss [33] is then
@@
where $U=\pi E_{0}^{2}\sigma_{0}^{2}t_{\text{int}}/2$ is the pulse energy [55] and $\sigma_{0}$ its focused spot size. This result shows that the energy loss in the classical regime depends on the fluence of the laser pulse, i.e., it is $\propto U/\sigma_{0}^{2}$, but is independent of the interaction time $t_{\text{int}}$ [56]. For a SFG pulse, the finite Rayleigh range $z_{R}=\omega_{0}\sigma_{0}^{2}/2$ can be taken into account by substituting $\xi_{0}\rightarrow\xi_{0}/\sqrt{1+z^{2}/z_{R}^{2}}=\xi_{0}/\sqrt{1+t^{2}/z_{R}^{2}}$ in the expression for $\chi$, such that $\chi_{\text{SFG}}(t)\approx 2\mathcal{E}(t)\omega_{0}\xi_{0}|\sin(2\omega_{0}t)|/m^{2}\sqrt{1+t^{2}/z_{R}^{2}}$. As a result, $\kappa_{\text{C,SFG}}=\kappa_{\text{C,FF}}\arctan(\rho)/\rho$, with $\rho\equiv t_{\text{int}}/2z_{R}$ [33]. Thus, for a fixed laser energy, a FF pulse and a short SFG pulse with $\rho\ll 1$ induce the same energy loss in the classical regime.
In the quantum regime ($\chi\gtrsim 1$), the emission rate [Eq. (1)] has a weaker dependence on the field amplitude than in the classical regime. This suggests that, in the quantum regime, a long FF pulse can induce more energy loss than a short SFG pulse with the same energy. This reasoning can be verified analytically for FF pulses in the limit $\chi\gg 1$, where $d\mathcal{E}/dt\propto-\chi^{2/3}$ and
@@
Thus, for a fixed pulse energy $U$, higher energy losses can be achieved by extending the interaction time $t_{\text{int}}$.
Figure 1 confirms the above expectations by showing the relative energy loss as a function of the interaction time for electrons colliding head-on with either FF pulses (blue solid lines) or SFG pulses (dashed orange lines) with energies of (a) 1 J and (b) 10 J. The solid and dashed lines were obtained by numerically integrating Eq. (1), with $\chi$ equal to $\chi_{\text{FF}}$ or $\chi_{\text{SFG}}$, respectively. In both cases, the maximum value of the quantum nonlinearity parameter is $\chi_{0}=2\mathcal{E}_{0}\omega_{0}\xi_{0}/m^{2}\approx 0.1\,\xi_{0}\mathcal{E}_{0}(10\;\text{GeV})$. For $U=$ 1 J, the maximum energy loss achieved by a FF pulse is approximately $30\%$ higher than that of an SFG pulse regardless of the initial electron energy. For $U=$ 10 J and $\mathcal{E}_{0}=$ 0.5 GeV, the maximum energy loss achieved by a FF pulse is $12\%$ higher. In this case and for larger electron energies, the enhanced energy loss in the FF pulses is reduced because the electron loses almost all of its energy whether it interacts with a FF or SFG pulse (see also the inset of Fig. 2a discussed below).
Figure 1 also shows that the energy loss in SFG pulses is suppressed for either very short pulses ($\rho\ll 1$) or very long pulses ($\rho\gg 1$). Shorter duration SFG pulses have higher peak intensities, but, as discussed above, the emission rate has a weaker scaling with intensity than it does with duration in the quantum regime [57]. With long duration SFG pulses, the electron spends a large fraction of the interaction time outside of the confocal region where the intensity is lower. This contrasts the FF pulses where the electron spends the entire interaction near the focus where the intensity is high. As a result, the energy loss increases monotonically with the interaction time until it approaches the classical limit [green dotted lines evaluated using Eq. (3)]: for fixed pulse energy, $\chi_{0}\propto t_{\text{int}}^{-1/2}$. The energy loss in the FF and SFG pulses coincide at short interaction times because the two pulses have the same amplitude and the finite Rayleigh range of the SFG pulse does not play a role ($\rho\ll 1$).
The results presented in Fig. 1 were verified by evolving electrons in the full FF and SFG fields with a Monte Carlo radiation emission that takes into account the correlations in multiple photon emission (see Section I of SM [58]). Notwithstanding, some remarks are in order to address the validity of the LCFA. The LCFA overestimates the exact QED emission spectrum for sufficiently low-energy photons [59, 60, 61]. The key parameter is [59]
@@
where $\omega$ is the energy of the emitted photon and $\xi$ and $\chi$ are the classical and quantum nonlinearity parameters at the instant of emission. The condition $\eta_{\text{LCFA}}^{2/3}\ll 1$ ensures that corrections to the LCFA are small [60]. Equation (1) is derived by integrating the emitted photon spectrum over the photon energy. For $\chi\ll 1$, the largest contribution to the integral comes from energies $\omega\lesssim\chi\mathcal{E}\ll\mathcal{E}$; for $\chi\sim 1$, the largest contribution comes from energies $\omega\lesssim\mathcal{E}$. As a result, the validity of the LCFA is determined by the low-energy photons and the condition $\eta_{\text{LCFA}}^{2/3}\sim 1/\xi^{2}\ll 1$. The vertical line in Fig. 1 corresponds to the choice $\xi_{0}=5$ [62] so that $\eta_{\text{LCFA}}^{2/3}=1/\xi_{0}^{2}=0.04$. Note that the expression for the classical energy loss does not rely on the validity of the LCFA, which explains why in the FF case the curves tend to the classical result for longer and longer interaction times ($\chi_{0}\propto t_{\text{int}}^{-1/2}$ for fixed energy).
Figure 2: Relative energy loss $\zeta$ as a function of the initial electron energy for electrons colliding head-on and on-axis with equal-energy FF and SFG pulses. In (a), the pulses have an energy $U=$ 1 J, 10 J, or 100 J and a spot size $\sigma_{0}=$1.5 $\mu$m. In (b), $U=$ 10 J and the spot size is $\sigma_{0}=$ 1.5 $\mu$m or 4 $\mu$m. The color scale shows the maximum quantum nonlinearity parameter. Above $\mathcal{E}\approx$ 3 GeV, both the FF and SFG are in the quantum regime. The inset shows the relative improvement in the energy loss afforded by the FF pulses. At $U=$ 1 J, the FF results in ${\sim}30\%$ more loss than the SFG. At $U=$ 10 J, both the FF and SFG begin to approach $100\%$ loss, which limits the relative improvement achievable with a FF pulse.
@@
To further illustrate the advantage of FF pulses in the quantum regime of radiation reaction, Fig. 2 compares the maximum energy loss induced by FF pulses (the longest $t_{\mathrm{int}}$ satisfying $\eta_{\text{LCFA}}^{2/3}\sim 1/\xi_{0}^{2}<0.04$) to that of SFG pulses (the maxima of the orange dashed curves in Fig. 1). For almost all initial electron energies $\mathcal{E}_{0}$, the FF pulses (solid lines) result in a larger loss of electron energy than the SFG pulses (dashed lines). However, as the pulse energy increases, the electron energy loss for both the FF and SFG pulses approaches unity (Fig. 2a), which limits the relative advantage of the FF pulses (Fig. 2a inset). In addition, both pulses achieve higher energy losses with smaller spot sizes (Fig. 2b) [see also Eq. (3) for the classical regime]. Thus, defocusing an SFG pulse to increase the interaction time would not improve its results.

### v2

@@ Abstract
The emission of a photon by an electron in an intense laser field is one of the most fundamental processes in electrodynamics and underlies the many applications that utilize high-energy photon beams. This process is typically studied for electrons colliding head-on with a stationary-focus laser pulse. Here, we show that the energy lost by electrons in the quantum regime and the yield of emitted photons can be substantially increased by replacing a stationary-focus pulse with an equal-energy flying-focus pulse whose focus co-propagates with the electrons. These advantages of the flying focus result from the energy loss and the photon yield scaling more favorably with the interaction time than the laser intensity in the quantum regime, with the latter also holding in the classical regime. Monte Carlo simulations of electrons colliding with equal-energy stationary and flying-focus laser pulses demonstrate these advantages.
I Introduction
@@
Modern high-power lasers enable direct probing of strong-field quantum electrodynamics with optical light [E144:1996enr, Burke:1997ew, Cole:2017zca, Poder:2017dpw, mirzaie2024, los2024observation]. Strong-field quantum electrodynamics (SFQED) refers to processes that occur in electromagnetic fields with amplitudes of the order of the Schwinger critical field $F_{\text{cr}}=m^{2}/|e|=1.3\times 10^{16}$ V/cm = $4.4\times 10^{13}$ G [DiPiazza:2011tq, Burton:2014wsa, Gonoskov:2021hwf, Fedotov:2022ely]. Here, $e<0$ is the electron charge, $m$ its mass, and units $\hbar=c=\epsilon_{0}=1$ are employed throughout. The corresponding intensity is $I_{\text{cr}}=4.6\times 10^{29}$ W/cm2. To date, the highest intensity achieved by a high-power laser in the laboratory frame is about six orders of magnitude below this value [Yoon:2021ony]. Nevertheless, the peak intensity experienced by an ultrarelativistic electron in its rest frame can readily exceed $I_{\text{cr}}$. The key parameter that determines the strength of a SFQED process is $\chi=\sqrt{-(F^{\mu\nu}p_{\nu})^{2}}/mF_{\text{cr}}$, where $p_{\nu}$ is the electron four-momentum, $F^{\mu\nu}$ is the electromagnetic tensor of the laser field, and the metric is $\eta^{\mu\nu}=\text{diag}(+1,-1,-1,-1)$. Because $\chi$ is a Lorentz-invariant quantity, it is directly proportional to the electric field experienced by an electron in its rest frame. As a result, a value of $\chi\gtrsim 1$ can be achieved experimentally by colliding available beams of multi-GeV electrons [Gonsalves:2019wnc, Abramowicz_2019, Meuren_2020, Picksley:2024cdd] with existing high-power laser pulses.
@@
Space-time structured laser pulses, such as the flying focus (FF), offer a new paradigm for probing SFQED processes. FF pulses feature a programmable velocity focus that moves independently of the group velocity [Sainte-Marie_2017, Froula_2018, Turnbull_2018]. The moving intensity peak formed by the FF can travel distances much longer than a Rayleigh range while maintaining a near-constant spatiotemporal profile [Froula_2018, Palastro_2018, Turnbull_2018, Howard_2019, simpson2020nonlinear, Ramsey_2020, Palastro:2020gcl, Jolly:2020, simpson2022spatiotemporal, Ramsey_2022, pigeon2024ultrabroadband, Formanek:2023zia, ramsey2023exact, gong2024laser]. Experimental configurations and conceptual proposals for producing FF pulses have used chromatic optics [Froula_2018, Turnbull_2018, Jolly:2020], axiparabola-echelon pairs [Palastro:2020gcl, smartsev2019axiparabola, oubrerie2022axiparabola, pigeon2024ultrabroadband], and nonlinear optical processes [simpson2020nonlinear, simpson2022spatiotemporal]. In the context of high-field physics, a FF pulse with an intensity peak that moves at the speed of light in the opposite direction of its phase fronts allows for extended interaction times with high-energy particles that collide head-on with the phase fronts. This configuration was first suggested to accentuate signatures of the transverse-formation-length in the emission of radiation [DiPiazza:2020wxp]. The same configuration was later proposed in Refs. [Formanek:2021bpw, Formanek:2023mkx] to facilitate the detection of classical radiation-reaction and vacuum birefringence. The observed effect in both cases, i.e., the classical energy loss of the electron beam and polarization rotation of an x-ray probe beam, scaled with the energy of the laser pulse [DiPiazza:2011tq, Burton:2014wsa, Gonoskov:2021hwf, Fedotov:2022ely]. Thus, the extended interaction time afforded by the FF enabled the use of much lower and more controllable powers than ultrashort, stationary-focus Gaussian (SFG) pulses. The magnitude of the effect produced by the FF and SFG was, however, the same.
@@
In this letter, we show that a FF pulse can significantly enhance observable effects of quantum radiation reaction when compared to a SFG with the same energy. More specifically, FF pulses result in a greater loss of electron energy and a larger yield of photons in the 1–20 MeV range than SFG pulses. This is because the electron energy loss scales more favorably with the interaction time than the laser intensity in the quantum regime ($\chi\gtrsim 1$), while the photon yield scales more favorably with the interaction time in both the classical and quantum regimes. A larger yield of photons in the 1–20 MeV range could impact several applications, including radiosurgery [girolami1996photon, weeks1997compton], photo-transmutation for treatment of long-lived nuclear waste and production of medical isotopes [wang2017transmutation, irani2014gamma, luo2016production], and investigations into the structure of materials using nuclear resonance fluorescence [kneissl1996investigation, albert2011design]. While many methods have been proposed to increase the yield of high-energy photons (see e.g., [powers2014quasi, vyskovcil2020inverse, luedtke2021creating, morris2021highly, hadjisolomou2023gamma, yu2024compact]), here, we show something qualitatively different: how space-time structured light can significantly enhance the emission probability of the underlying fundamental process.
II Radiation energy loss
@@
The basic SFQED process describing radiation by an ultrarelativistic electron beam in the field of a laser pulse is nonlinear Compton scattering. Here, this process will be investigated within the locally-constant field approximation (LCFA) [DiPiazza:2011tq, Burton:2014wsa, Gonoskov:2021hwf, Fedotov:2022ely]. The LCFA is a useful approximation to study SFQED phenomena because it allows one to compute the probability of a process in an arbitrary electromagnetic field using a known expression for the probability in a constant crossed field. A constant crossed field is a space and time-independent electromagnetic field $(\bm{E},\bm{B})$ that satisfies $|\bm{E}|=|\bm{B}|$ and $\bm{E}\cdot\bm{B}=0$ [Ritus:1985vta, Reiss:1962nhe, Baier:1998vh]. The LCFA is valid provided that: (a) ultrarelativistic charges move fast enough to experience an arbitrary electromagnetic field as an approximate crossed field in their rest frame [Jackson_b_1975] and (b) the formation length of a process like nonlinear Compton scattering is much smaller than the wavelength of the laser pulse $\lambda_{0}=2\pi/\omega_{0}$, where $\omega_{0}$ is the central angular frequency of the pulse. This latter condition is typically satisfied when the dimensionless field strength $\xi_{0}=|e|E_{0}/m\omega_{0}\gg 1$, where $E_{0}$ is the electric field amplitude of the pulse [Ritus:1985vta, Baier:1998vh].
@@
Within the LCFA, the average energy $\mathcal{E}_{\gamma}$ radiated by an electron per unit time in a background electromagnetic field through emission of a single photon can be calculated using the Baier-Katkov interpolation formula [Baier:1998vh]
@@
where $\alpha\approx 1/137$ is the fine-structure constant and $\chi$ is evaluated using the local values of $p_{\nu}$ and $F^{\mu\nu}$. The interpolation has an accuracy better than 2% for all values of $\chi$ [Baier:1998vh]. In the regime under consideration where the electron is ultrarelativistic ($\xi_{0}\gg 1$ and $\chi\sim 1$), the likelihood of multiple photon emission within a formation length is suppressed. Further, the work done by a laser pulse within a formation length is negligible compared to the energy lost to photon emission. Thus, the quantity $d\mathcal{E}_{\gamma}/dt$ is approximately equal to the average energy lost by an electron per unit time, i.e., $d\mathcal{E}/dt=-d\mathcal{E}_{\gamma}/dt$, where $\mathcal{E}=\mathcal{E}(t)$ is the electron energy at time $t$ [Glauber_1951, Di_Piazza_2010]. Equation (1) can then be used to calculate the relative electron energy loss $\zeta$ defined as
@@
where $\mathcal{E}_{0}$ is the initial energy of an electron and $\mathcal{E}_{F}$ its energy after interacting with the laser pulse. Equation 1 applies to a single electron but can be interpreted as the continuous, average energy loss per-electron for an ensemble of electrons sharing the same initial condition. This is verified in Section I of the Supplemental Material (SM) [supplemental] by comparing the results of Eq. (1) to a stochastic quantum mechanical approach (see also Ref. [niel2018quantum]).
[…]
The advantage of using a FF pulse in place of a SFG pulse with the same energy will be demonstrated by numerically integrating Eq. (1). An analytical treatment is also possible by making a few simplifying assumptions: (1) $\mathcal{E}\gg m\xi_{0}$, which guarantees that, for an electron initially counterpropagating with respect to the laser pulse, the transverse electron motion and pulse structure can be ignored; (2) The focal point of the laser pulse is located at $x=y=0$ and the electron moves along the negative $z$ axis at an ultrarelativistic velocity such that $z(t)\approx z(0)-t$; (3) The temporal profile of the pulse has a square shape equal to unity with a duration $2t_{\text{int}}$, which intersects the electron from $z=t_{\text{int}}/2$ to $z=-t_{\text{int}}/2$.
In the classical regime ($\chi\ll 1$), Eq. (1) can be approximated as $d\mathcal{E}/dt\approx-2\alpha m^{2}\chi^{2}/3$. When interacting with the FF pulse, the electron experiences a field with a near-constant amplitude $\xi_{0}$ such that $\chi_{\text{FF}}(t)\approx 2\mathcal{E}(t)\omega_{0}\xi_{0}|\sin(2\omega_{0}t)|/m^{2}$. The relative energy loss [Formanek:2021bpw] is then
@@
where $U=\pi E_{0}^{2}\sigma_{0}^{2}t_{\text{int}}/2$ is the pulse energy [Esarey:1993zz] and $\sigma_{0}$ its focal spot size. This result shows that the energy loss in the classical regime depends on the fluence of the laser pulse, i.e., it is $\propto U/\sigma_{0}^{2}$, but is independent of the interaction time $t_{\text{int}}$  [^1] Note that subleading oscillatory terms were ignored.. For a SFG pulse, the finite Rayleigh range $z_{R}=\omega_{0}\sigma_{0}^{2}/2$ can be taken into account by substituting $\xi_{0}\rightarrow\xi_{0}/\sqrt{1+z^{2}/z_{R}^{2}}=\xi_{0}/\sqrt{1+t^{2}/z_{R}^{2}}$ in the expression for $\chi$, such that $\chi_{\text{SFG}}(t)\approx 2\mathcal{E}(t)\omega_{0}\xi_{0}|\sin(2\omega_{0}t)|/m^{2}\sqrt{1+t^{2}/z_{R}^{2}}$. As a result, $\kappa_{\text{C,SFG}}=\kappa_{\text{C,FF}}\arctan(\rho)/\rho$, with $\rho\equiv t_{\text{int}}/2z_{R}$ [Formanek:2021bpw]. Thus, for a fixed laser energy, a FF pulse and a short SFG pulse with $\rho\ll 1$ induce the same energy loss in the classical regime.
In the quantum regime ($\chi\gtrsim 1$), the scalings change because the emission rate [Eq. (1)] has a weaker dependence on the field amplitude than in the classical regime. This suggests that, in the quantum regime, a long FF pulse can induce more energy loss than a short SFG pulse with the same energy. This reasoning can be verified analytically for FF pulses in the limit $\chi\gg 1$, where $d\mathcal{E}/dt\propto-\chi^{2/3}$ and
@@
Thus, for a fixed pulse energy $U$, higher energy losses can be achieved by extending the interaction time.
Figure 1 confirms the above expectations by showing the relative energy loss as a function of the interaction time for electrons colliding head-on with either a FF pulse (blue solid lines) or a SFG pulse (dashed orange lines) with energies of (a) 1 J or (b) 10 J. The solid and dashed lines were obtained by numerically integrating Eq. (1), with $\chi$ equal to $\chi_{\text{FF}}$ or $\chi_{\text{SFG}}$, respectively. In both cases, the maximum value of the quantum nonlinearity parameter is $\chi_{0}=2\mathcal{E}_{0}\omega_{0}\xi_{0}/m^{2}\approx 0.1\,\xi_{0}\mathcal{E}_{0}(10\;\text{GeV})$. For $U=$ 1 J, the maximum energy loss achieved by a FF pulse is approximately $30\%$ higher than that of an SFG pulse regardless of the initial electron energy. For $U=$ 10 J and $\mathcal{E}_{0}=$ 0.5 GeV, the maximum energy loss achieved by a FF pulse is $12\%$ higher. In this case and for larger electron energies, the enhanced energy loss in the FF pulses is reduced because the electron loses almost all of its energy whether it interacts with a FF or SFG pulse (see also the inset of Fig. 2a discussed below).
Figure 1 also shows that the energy loss in SFG pulses is suppressed for either very short pulses ($\rho\ll 1$) or very long pulses ($\rho\gg 1$). Shorter duration SFG pulses have higher peak intensities, but, as discussed above, the emission rate has a weaker scaling with intensity than it does with duration in the quantum regime [niel2018quantum]. With long duration SFG pulses, the electron spends a large fraction of the interaction time outside of the confocal region where the intensity is lower. This contrasts with FF pulses where the electron spends the entire interaction near the focus where the intensity is high. As a result, the energy loss increases monotonically with the interaction time until it approaches the classical limit [green dotted lines evaluated using Eq. (3)]: for fixed pulse energy, $\chi_{0}\propto t_{\text{int}}^{-1/2}$. The energy loss in the FF and SFG pulses coincide at short interaction times because the two pulses have the same amplitude and the finite Rayleigh range of the SFG pulse does not play a role ($\rho\ll 1$).
The results presented in Fig. 1 were verified by evolving electrons in the full FF and SFG fields with a Monte Carlo radiation emission calculation that takes into account the stochastic nature of multiple photon emission (see Section I of SM [supplemental]). Notwithstanding, some remarks are in order to address the validity of the LCFA. The LCFA overestimates the amplitude of the exact QED emission spectrum for sufficiently low photon energies [Di_Piazza_2018, Di_Piazza_2019, Ilderton_2019_b]. The key parameter is [Di_Piazza_2018]
@@
where $\omega$ is the energy of the emitted photon and $\xi$ and $\chi$ are the classical and quantum nonlinearity parameters at the instant of emission. The condition $\eta_{\text{LCFA}}^{2/3}\ll 1$ ensures that corrections to the LCFA are small [Di_Piazza_2019]. Equation (1) is derived by integrating the emitted photon spectrum over the photon energy. For $\chi\ll 1$, the largest contribution to the integral comes from energies $\omega\lesssim\chi\mathcal{E}\ll\mathcal{E}$; for $\chi\sim 1$, the largest contribution comes from energies $\omega\lesssim\mathcal{E}$. As a result, the validity of the LCFA is determined by the low-energy photons and the condition $\eta_{\text{LCFA}}^{2/3}\sim 1/\xi^{2}\ll 1$. The vertical line in Fig. 1 corresponds to the choice $\xi_{0}=5$ [Blackburn:2018sfn] so that $\eta_{\text{LCFA}}^{2/3}=1/\xi_{0}^{2}=0.04$. Note that the expression for the classical energy loss does not rely on the validity of the LCFA, which explains why in the FF case the curves tend to the classical result as the interaction time increases ($\chi_{0}\propto t_{\text{int}}^{-1/2}$ for fixed energy).
[…]
@@
To further illustrate the advantage of FF pulses in the quantum regime of radiation reaction, Fig. 2 compares the maximum energy loss induced by FF pulses (with the longest $t_{\mathrm{int}}$ satisfying $\eta_{\text{LCFA}}^{2/3}\sim 1/\xi_{0}^{2}<0.04$) to that of SFG pulses (the maxima of the orange dashed curves in Fig. 1). For almost all initial electron energies $\mathcal{E}_{0}$, the FF pulses (solid lines) result in a larger loss of electron energy than the SFG pulses (dashed lines). However, as the pulse energy increases, the electron energy loss for both the FF and SFG pulses approaches unity (Fig. 2a), which limits the relative advantage of the FF pulses (Fig. 2a inset). In addition, both pulses achieve higher energy losses with smaller spot sizes (Fig. 2b) [see also Eq. (3) for the classical regime]. Thus, defocusing an SFG pulse to increase the interaction time would not improve its results.
Figure 3: Number of photons emitted per electron at $\chi>0.1$ from the on-axis collision of $\mathcal{E}_{0}=10$ GeV electrons with 1 J FF and SFG laser pulses. The inset shows the number of photons emitted as a function of energy and $\eta_{\text{LCFA}}^{2/3}$ for the FF pulse. The numbers in white indicate what percentage of photons in the given range were emitted at $\chi>0.1$.
II.1 Photon yield